x/3-1/2=x+1/3 how do you find the answer

Answer by stanbon(75887)

x/3-1/2=x+1/3
--------
Multiply thru by 6 to get:
2x - 3 = 6x + 2
----
4x = -5
x = -5/4
-----------------
Cheers,
Stan H.